OpenStudy (anonymous):

The polynomial x^2 + 16x + 100 can be written as (x + a)^2 + b where a = ? and b = ?

OpenStudy (anonymous):

\[add and subtract \left( \frac{ co-efficient of x }{ 2 } \right)^{2} i.e.,\left( \frac{ 16 }{ 2 } \right)^{2},or 64\]